# Enhancing capabilities of particle image/tracking velocimetry with physics-informed neural networks

### v2

@@ IV Results
[…]
$\Delta_{Q}=\left\langle\dfrac{||Q-Q_{DNS}||_{2}}{||Q_{DNS}||_{2}}\right\rangle_{T},$ (10)
where $Q$ denotes the physical quantity of interest predicted by the PINN, and $Q_{DNS}$ is its exact value. Table 1 shows the results for velocity, pressure and external force, as well as the relative errors in determining the bottom friction $\alpha$ and fluid viscosity $\nu$. The developed method consistently reconstructs the velocity field in the observation region with an accuracy of about $0.2\%$, while the model parameters $\alpha$ and $\nu$ are determined with an error of several percent. Relative errors in predicting the pressure field and external force are larger, with greater dispersion and an obvious correlation between them. The average for both is about $8\%$, and the standard deviation is close to $5\%$. This behavior is due to the lack of measurements for these quantities, so the model extracts them solely from the analysis of physical laws.
Table 1: Relative $L_{2}$ errors in predicting velocity, pressure, external force, bottom friction and fluid viscosity (with clean data).
model $\Delta_{\bm{v}},\%$ $\Delta_{p-\bar{p}},\%$ $\Delta_{\bm{f}},\%$ $\Delta_{\alpha},\%$ $\Delta_{\nu},\%$
$A_{0}$ $0.22$ $5.52$ $6.09$ $3.66$ $1.42$
$A_{1}$ $0.21$ $4.21$ $4.02$ $4.27$ $1.56$
$A_{2}$ $0.20$ $3.11$ $4.10$ $2.83$ $1.41$
$A_{3}$ $0.23$ $17.34$ $16.58$ $4.36$ $1.50$
$A_{4}$ $0.20$ $2.13$ $2.28$ $3.22$ $1.26$
$A_{5}$ $0.24$ $8.34$ $8.39$ $3.84$ $1.76$
$A_{6}$ $0.22$ $10.46$ $10.17$ $4.02$ $1.47$
$A_{7}$ $0.21$ $17.37$ $16.59$ $4.21$ $1.53$
$A_{8}$ $0.21$ $3.85$ $3.66$ $4.07$ $1.65$
$A_{9}$ $0.23$ $7.48$ $7.25$ $4.10$ $1.59$
@@
Table 2: Dependence of mean values and standard deviations for prediction errors on the noise level.
noise $\overline{\Delta}_{\bm{v}},\%$ $\overline{\Delta}_{p-\bar{p}},\%$ $\overline{\Delta}_{\bm{f}},\%$ $\overline{\Delta}_{\alpha},\%$ $\overline{\Delta}_{\nu},\%$
$0\%$ $0.22\pm 0.01$ $8.0\pm 5.6$ $7.9\pm 5.2$ $3.9\pm 0.5$ $1.5\pm 0.1$
$0.5\%$ $0.28\pm 0.01$ $8.9\pm 5.4$ $8.6\pm 5.1$ $3.1\pm 0.9$ $1.2\pm 0.2$
$1\%$ $0.50\pm 0.03$ $7.8\pm 5.9$ $7.6\pm 5.6$ $2.2\pm 1.8$ $0.7\pm 0.5$
$2\%$ $2.39\pm 0.06$ $12.9\pm 5.1$ $12.2\pm 5.3$ $41.4\pm 4.4$ $12.3\pm 0.9$
$5\%$ $5.84\pm 0.28$ $19.1\pm 5.5$ $17.9\pm 6.0$ $188\pm 26$ $56.1\pm 2.0$
@@
Next, we repeat the training, but now the models receive noisy measurement data $\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})$. Table 2 reports the dependence of the average values and standard deviations for prediction errors on the noise level, obtained by averaging over ten models in each case. The developed algorithm is robust to small noise ($\leq 1\%$) in the input data. Moreover, the error for the reconstructed velocity field in this case is less than the noise level. Thus, PINN is able to correct for small noise based on the knowledge that the velocity field must satisfy the Navier-Stokes equation and incompressibility condition. As the noise level increases ($\geq 2\%$), the performance of the trained models gradually decreases. Since for a turbulent flow the dissipative terms make a small contribution to the Navier-Stokes equation, the model quickly loses the ability to predict the bottom friction $\alpha$ and fluid viscosity $\nu$. At the same time, the accuracy of the velocity field determination remains comparable to the noise level. Errors in predicting the pressure field and external force are larger and amount to tens of percent.
To characterize the statistical properties of reconstructed dense velocity, pressure and force fields, we calculate the corresponding probability density functions (PDFs); see Fig. 4. To illustrate, we used trained models with pressure and force errors close to the mean values presented in Table 2. The velocity PDFs are almost identical to the DNS data even for PINN models based on noisy input data, confirming the low $L_{2}$ error between them. The pressure field is reconstructed worse, and the main difference from the DNS data lies in the typical values, and not in the tails of the distribution function. The force PDF corresponding to expression (2) can be calculated analytically
$\rho(f)=\dfrac{2}{\pi f_{0}}\left(1-\dfrac{f^{2}}{f_{0}^{2}}\right)^{-1/2},$ (11)
where $f=(f_{x}^{2}+f_{y}^{2})^{1/2}\leq f_{0}$, and it diverges at $f\to f_{0}$. The largest disparities between the reconstructed PDFs and the DNS data are concentrated around this point.
@@
In Fig. 5, we compare the energy spectra of the reference flow and PINN reconstructions, which are commonly used to characterize the statistical properties of a velocity field. Since the observation region is not periodic, when constructing the spectra we applied a Hann window function. For this reason, the DNS spectrum is slightly different from that presented earlier in Fig. 2b. As expected from the previous analysis, the overall errors are small. They are mainly concentrated on small scales with $k\gg k_{f}$ and gradually rise with increasing noise level in the input data.
V Conclusion
@@
We showed that PINN can combine measurement data with an accurate physical model to extract additional information from ordinary flow observations. The developed technique can be built on top of current PIV/PTV workflows, potentially improving their functionality. In particular, we investigated a forced two-dimensional turbulent fluid flow and demonstrated that the proposed method can reconstruct the dense velocity and pressure fields, infer the driving force, and determine the unknown fluid viscosity and bottom friction coefficient using only sparse and noisy velocity measurements.
@@
The suggested technique trains two independent NNs simultaneously, one to estimate the velocity and pressure field and the other to approximate the external forcing. This enables us to use different designs in NNs, which makes it possible to take into account the physical features of the considered system. In our case, this allowed us to take into account the stationarity of the external forcing implemented in the physical model. Training of NNs involves minimization of the loss function that penalizes deviations from measured data and violations of the Navier-Stokes equation and incompressibility conditions. While our intention was to select the optimal NN design and hyperparameter configuration, a thorough examination of this issue was outside the scope of this work. We anticipate that the introduction of more advanced NN architectures Zhao, Ding, and Prakash (2023); Wang et al. (2024); Liu et al. (2024); Buzaev et al. (2023) can further improve the developed method.
@@
The reconstructed flows were compared to the reference flow in terms of $L_{2}$ errors, PDFs of the reconstructed fields, and energy spectra. Overall, the results show good accuracy for super-resolution and inference applications and demonstrate moderate robustness to noise in the input data. At low noise levels, the algorithm can correct errors in the velocity field measurements based on information from the physical model. The developed technique can be useful for analyzing flow observations when the physical model of the phenomenon is well constructed.

### v3

@@ V Results
$\delta_{Q}=\left\langle\dfrac{||\hat{Q}-Q||_{2}}{||Q||_{2}}\right\rangle_{T},$ (8)
where $\hat{Q}$ denotes the physical quantity of interest predicted by PINN, and $Q$ is its exact value from DNS. We will look at how well velocity, pressure, external force, bottom friction, and fluid viscosity are reconstructed. However, as discussed in Section II, the pressure and external force cannot be determined unambiguously from velocity field measurements. Therefore, to determine the accuracy of reconstructions of these fields, we will analyze errors in determining the pressure Laplacian $\nabla^{2}p$ and the external force curl $\phi=\partial_{x}f_{y}-\partial_{y}f_{x}$. Furthermore, to test the statistical properties of the predicted velocity, we will compare the energy spectra of the reference and reconstructed flows.
@@
Table 1: Dependence of mean values and standard deviations for prediction errors on the noise level $\varepsilon$.
$\varepsilon$ $\delta_{\bm{v}},\%$ $\delta_{\nabla^{2}p},\%$ $\delta_{\phi},\%$ $\delta_{\alpha},\%$ $\delta_{\nu},\%$ $\delta_{p-\langle p\rangle_{T}},\%$
00 $0.22\pm 0.01$ $4.0\pm 0.2$ $1.6\pm 0.3$ $3.9\pm 0.5$ $1.5\pm 0.1$ $0.58\pm 0.03$
$0.005$ $0.28\pm 0.01$ $4.7\pm 0.1$ $1.7\pm 0.4$ $3.1\pm 0.9$ $1.2\pm 0.2$ $0.86\pm 0.03$
$0.01$ $0.50\pm 0.03$ $10.1\pm 1.8$ $2.0\pm 0.4$ $2.2\pm 1.8$ $0.7\pm 0.5$ $1.7\pm 0.1$
$0.02$ $2.39\pm 0.06$ $157\pm 9$ $5.4\pm 0.8$ $41.4\pm 4.4$ $12.3\pm 0.9$ $7.7\pm 0.3$
$0.05$ $5.84\pm 0.28$ $478\pm 43$ $13.0\pm 2.7$ $188\pm 26$ $56.1\pm 2.0$ $17.3\pm 1.0$
V.1 Impact of input data density
First, we trained PINN models based on noise-free input data. We set a fixed number $N_{eq}=9\times 10^{4}$ of collocation points and vary the number $N_{data}$ of data points in the range from $3\times 10^{2}$ to $9\times 10^{4}$ to study the dependence of prediction accuracy on input data density. Reconstructions employing input data with $N_{data}\geq 10^{4}$ are highly accurate. Fig. 3 compares DNS data with NN predictions in the observation zone at some instant of time for a PINN model trained with $N_{data}=3\times 10^{4}$ measurement points. The outcomes appear to be very close, with differences mostly localized on small scales.
Fig. 4 reports RRMSEs for determining the velocity, pressure Laplacian, external force curl, bottom friction, and fluid viscosity. As $N_{data}$ increases, RRMSEs first decrease rapidly in a power-law manner ($N_{data}\leq 10^{4}$), and then the dependence gradually begins to saturate ($N_{data}\geq 10^{4}$). Similar behavior was observed earlier for the simpler problem of reconstructing 2D Taylor’s decaying vortices Wang, Liu, and Wang (2022). The velocity reconstruction error is less than $0.5\%$ for $N_{data}\geq 10^{4}$, while for other quantities the errors are approximately an order of magnitude larger. This behavior is due to the lack of measurements for these quantities, so the model extracts them solely from the analysis of governing equations.
Figure 5: Energy spectra for DNS data and PINN models trained with different numbers $N_{data}$ of data points.
In Fig. 5, we compare the energy spectra of the reference flow and PINN reconstructions, which are commonly used to characterize the statistical properties of a velocity field. Since the observation region is not periodic, when constructing the spectra we applied a Hann window function. For this reason, the DNS spectrum is slightly different from that presented earlier in Fig. 2b. The statistical properties of the velocity field are well reproduced for $N_{data}\geq 10^{3}$. Any noticeable differences are observed only for more sparse data, and they are mainly localized at small scales with $k\gg k_{f}$.
V.2 Noisy velocity measurements
@@
Next, we fix $N_{data}=3\times 10^{4}$ and repeat the training, but now the models receive noisy measurement data. Table 1 reports the dependence of the average values and standard deviations for prediction errors on the noise level, obtained by averaging over ten models in each case. The developed algorithm is robust to small noise ($\varepsilon\leq 0.01$) in the input data. Moreover, the error for the reconstructed velocity in this case is less than the noise level. Thus, PINN is able to correct for small noise based on the knowledge that the velocity field must satisfy the Navier-Stokes equation and incompressibility condition.
As the noise level increases ($\varepsilon\geq 0.02$), the performance of the trained models gradually decreases. Since for a turbulent flow the dissipative terms make a relatively small contribution to the Navier-Stokes equation, the model quickly loses the ability to predict the bottom friction $\alpha$ and fluid viscosity $\nu$. At the same time, the accuracy of the velocity field and external force determination remains comparable to the noise level.
@@
Errors in predicting the pressure Laplacian take on gigantic values, but in fact this is due to the peculiarities of the introduced metric, and the pressure field itself is reconstructed quite well. The presence of noise in the training data leads to small-scale errors, which are amplified when calculating the second-order spatial derivatives. A similar statement applies to the interpretation of errors in determining the external force curl, but to a lesser extent, since in this case only first-order derivatives are calculated. Fig. 6 compares the energy spectra for the reference and reconstructed flows, and supports the statement that there are small-scale errors in the predicted velocity.
To evaluate the accuracy of pressure reconstruction in a more direct way, we consider the modified pressure $p-\langle p\rangle_{T}$. Subtracting the mean value over time allows one to eliminate the ambiguity in determining pressure associated with an arbitrary harmonic function $h(\bm{r})$, see details in Section II. The modified pressure is determined up to an arbitrary constant, so when comparing these fields we also subtract the spatial average values from them. The results are illustrated in Fig. 7, where we compare the modified pressures at some moment of time for DNS data and a PINN model trained on velocity measurements with a noise level of $\varepsilon=0.05$. Despite the noticeable differences, the modified pressures are generally similar to each other. The corresponding RRMSEs for all noise levels are shown in the last column of Table 1 and they are significantly smaller than the errors in the pressure Laplacian reconstruction.
Figure 7: Comparison of modified pressure $p-\langle p\rangle_{T}$ for DNS data and a PINN model trained on noisy input data with $\varepsilon=0.05$.
VI Conclusion
@@
We showed that PINN can combine dense in time, spatially sparse and probably noisy velocity measurements with an accurate physical model to extract additional information from ordinary flow observations. The developed technique can be built on top of current PIV/PTV workflows, potentially improving their functionality. In particular, we investigated a forced two-dimensional turbulent fluid flow and demonstrated that the proposed method can reconstruct the spatially dense velocity and pressure fields, infer the driving force, and determine the unknown fluid viscosity and bottom friction coefficient. The ability to reconstruct the full flow state from limited observations can also be used to compress experimental data.
@@
The suggested technique is based on training NNs by minimizing a loss function that penalizes deviations from measured data and violations of the governing equations. While our intention was to select the optimal NN design and hyperparameter configuration, a thorough examination of this issue was outside the scope of this work. We anticipate that the introduction of more advanced NN architectures Zhao, Ding, and Prakash (2023); Wang et al. (2024); Liu et al. (2024); Buzaev et al. (2023) can further improve PINN method.
@@
The reconstructed flows were compared to the reference flow in terms of relative root-mean-square errors and energy spectra. Overall, the results show good accuracy for super-resolution and inference applications and demonstrate moderate robustness to noise in the input data. At low noise levels, the algorithm can correct errors in the velocity measurements based on governing equations. The developed technique can be useful for analyzing flow observations when the physical model of the phenomenon is well constructed.
Although our study used time-dense velocity measurements, we would like to point out that the PINN method can be potentially extended and applied to time-sparse measurements as well. In this case, the interval between successive images of tracer particles is long enough and PIV/PTV methods are ineffective for determining instantaneous velocities. However, it is possible to complement the physical model with an equation describing the advection of tracer particles, and then train PINN on their experimental images. Examples of this strategy can be found in Ref. [Raissi, Yazdani, and Karniadakis, 2020], and instead of tracer particles, it is also possible to monitor the temperature field Cai et al. (2021b); Clark Di Leoni et al. (2023b). For turbulent systems, the distance between successive data frames cannot exceed the Lyapunov time, but how close one can get to this limit is an open question.
Finally, while the presented results look promising, the study did not address a number of issues that may be relevant for practical applications:
  • In many cases, the adopted physical model only approximately describes the system. For flows in thin fluid layers, equation (1) is only an approximation, and it breaks down for small-scale modes whose size is comparable to the fluid layer thickness. How sensitive is the PINN algorithm to modeling flaws?
  • The flow under consideration was spatially uniform. Will the algorithm work equally well for substantially non-uniform systems? What modifications would it require?
  • The procedure used to model the velocity measurements did not take into account all the features of experimental PIV/PTV measurements. It would be helpful to investigate the algorithm’s capabilities using experimental data collected under various settings. The first steps in this direction were done in Refs. [Wang, Liu, and Wang, 2022; Clark Di Leoni et al., 2023a; Cai, Gray, and Karniadakis, 2024].
  • How increasing the degree of flow non-linearity affect the performance of the algorithm? And how much data are required to achieve the desired reconstruction accuracy? Is it possible to approach the problem theoretically and find fundamental limitations on the required spatio-temporal data density?
We hope that future research will address these challenges and that the solutions will contribute to the further development of PINN.